Calculate the molar concentration of H+, HCO3-, and CO3-2 in a 0.200M Na2CO3 solution adjusted to pH 8.10.
Ka1 = 4.4 x 10^-7
Ka2 = 4.7 x 10^-11
1
Expert's answer
2020-04-18T07:00:48-0400

Ka1 = "{\frac {[H^+][HCO_3^-]} {[H_2CO_3]}}" = 4.4*10-7

Ka2 = "{\frac {[H^+][CO_3^{2-}]} {[HCO_3^-]}}" = 4.7*10-11

Na2CO3 -> 2Na+ + CO32-

Ka1*Ka2 = "{\frac {[H^+][HCO_3^-]} {[H_2CO_3]}}* {\frac {[H^+][CO_3^{2-}]} {[HCO_3^{-}]}}" ="{\frac {[H^+]^2[CO_3^{2-}]} {[H_2CO_3]}}=4.4*10^{-7}*4.7*10^{-11}=2.068*10^{-17}"

[H+] = 10^(-pH) = 10^(-8.10) = 7.943*10-9

"{\frac {[H^+]^2[CO_3^{2-}]} {[H_2CO_3]}}={\frac {(7.943*10^{-9})^2[CO_3^{2-}]} {[H_2CO_3]}}=2.068*10^{-17}"


"{\frac {[CO_3^{2-}]} {[H_2CO_3]}}= {\frac {2.068*10^{-17}} {6.309*10^{-17}} }=0.3278"

[CO32-] = x then [H2CO3] = 3.051x

the total amount of [CO32-], [HCO3-] and [H2CO3] is 0.2 M. Then [HCO3-] is 0.2 - x - 3.051x = 0.2-4.051x

Ka1 ="{\frac {[H^+][HCO_3^-]} {[H_2CO_3]}}={\frac {(7.943*10^{-9})(0.2-4.051x)} {3.051x}}=4.4*10^{-7}"

Solving the equation we get

x = 0.001156 M = 1.156 * 10-3 M = [CO32-]

0.2-4.051x = 0.1953 M = [HCO3-]

The answer:

[H+] = 7.943 * 10^-9 M

[HCO3-] = 1.953 * 10^-1 M

[CO32-] = 1.156 * 10^-3 M